# Movement-Prediction-Adjusted Naïve Forecast

### v7

@@ Abstract
This study introduces a movement-prediction-adjusted naïve forecast, which is the original naïve forecast with the addition of a weighted movement prediction term, in the context of forecasting time series that exhibit symmetric random walk properties. The weight of the movement term is determined by two parameters: one reflecting the directional accuracy and the other representing the mean absolute increment. The settings of the two parameters involve a trade-off: larger values may yield meaningful gains over the original naïve forecast, whereas smaller values often render the adjusted forecast more reliable. This trade-off can be managed by empirically setting the parameters using sliding windows on in-sample data. To statistically test the performance of the adjusted naïve forecast under different directional accuracy levels, we used four synthetic time series to simulate multiple forecast scenarios, assuming that for each directional accuracy level, diverse movement predictions were provided. The simulation results show that as the directional accuracy increases, the error of the adjusted naïve forecast decreases. In particular, the adjusted naïve forecast achieves statistically significant improvements over the original naïve forecast, even under a low directional accuracy of slightly above 0.50. This finding implies that the movement-prediction-adjusted naïve forecast can serve as a new optimal point forecast for time series with symmetric random walk characteristics if consistent movement prediction can be provided.
@@ 1 Introduction
Time series forecasting plays a crucial role in many decision-making processes. The widespread demand for accurate point forecasts across various domains has attracted considerable attention from researchers and practitioners. However, not all time series are equally predictable. Time series exhibiting random walk characteristics pose significant forecasting challenges (Hamilton,, 2020). Therefore, financial time series, known for these characteristics and their practical importance in financial markets, have been extensively studied as test cases for forecasting theories and methods. One prominent theory is the efficient market hypothesis (EMH), which suggests that asset prices incorporate all available information, rendering historical prices unpredictable (Fama,, 1970). For a random walk-like series, the naïve forecast, which assumes that the future value at the next time step is equal to the current value, is often regarded as the optimal point forecast.
@@
In support of the EMH, Moosa, (2013) reported that as the volatility of financial time series increases, the root mean square error (RMSE) of the forecasting models increases more quickly than that of the naïve forecast. Moreover, Moosa and Burns, (2014) argued that claims of outperforming the naïve forecast in terms of the RMSE are often misleading, frequently involving introduced dynamics without rigorous statistical validation. Despite attempts to use various predictors to break the assumed randomness of financial time series, existing forecasting methods, including statistical, machine learning, and deep learning approaches, typically cannot outperform naïve forecasts (Kilian and Taylor,, 2003; Moosa and Burns,, 2016; Thakkar and Chaudhari,, 2021; Petropoulos et al.,, 2022; Ellwanger and Snudden,, 2023; Hewamalage et al.,, 2023; Zeng et al.,, 2023; Beck et al.,, 2025).
@@
From the forecaster’s perspective, the unpredictability of random walk-like series could be attributed to incomplete knowledge of the underlying processes controlling the data, ignorance of the various contributing factors, and imperfections in analytical methods (Laplace,, 2012). In other words, the randomness of data is not inherent but rather apparent (Crutchfield and Feldman,, 2003). Although useful information may be contained in the input space, it can be overwhelmed by irrelevant data, making it difficult for current regression methods to exploit such information (John et al.,, 1994; Guyon and Elisseeff,, 2003; Verleysen and François,, 2005). Moreover, identifying the predictors on which the target variable truly depends is a complex task (Granger,, 1969; Stock and Watson,, 2002). However, these theories and empirical evidence reveal only the practical difficulty of finding exploitable patterns; they do not eliminate the possibility of discovering any predictability in the data.
@@
Given the challenge in the point forecasting of random walk-like series, many researchers have shifted their focus from point forecasting to movement prediction, which involves predicting whether the value is increasing or decreasing for the next time step. Consequently, the regression task was reduced to a binary classification task. For a pure random walk, the probability that future movement will increase or decrease is 0.5, and the mutual information between the predictors and the target is zero (Kraskov et al.,, 2004). If the mutual information is greater than zero, it indicates that there is some dependency between the predictors and the target (Cover and Thomas,, 2012). Practically, this means that the predictors reduce the uncertainty of the binary outcome. Although this does not guarantee high predictive accuracy, it suggests that, theoretically, one can build a model that performs better than random guessing (Ircio et al.,, 2020). In contrast, for point forecasting in a random walk-like series, the continuous nature of the target means that the small gain in mutual information only marginally reduces the overall uncertainty and does not sufficiently narrow down the precise numerical result (Beraha et al.,, 2019).
@@
Movement prediction is often considered easier than point forecasting in financial time series (Taylor,, 2008; Moosa,, 2013). Studies in the field of behavioral finance suggest that investor sentiment and psychology can lead to certain patterns (De Bondt and Thaler,, 1985; Lee et al.,, 1991; Baker and Wurgler,, 2007; Barberis et al.,, 1998), which could lead to mutual information between the sentiment predictors and the binary target being greater than zero and making future movements predictable. Therefore, sentiment indicators are often used to improve the directional accuracy in these tasks. For example, Weng et al., (2017) incorporated Google news counts and Wikipedia page views alongside historical data to predict the next day’s movement of AAPL stock via support vector machines (SVMs), achieving up to 0.80 directional accuracy. Ma et al., (2023) proposed a multisource aggregated classification (MAC) method for the prediction of stock price movement, incorporating the numerical features and market-driven news sentiments of target stocks, as well as the news sentiments of their related stocks, and the directional accuracy can reach as high as 0.70. Wang et al., (2025) used multimodal data that included news data to train a hybrid model consisting of a long short memory network (LSTM) and transformer for the prediction of stock movement, and the directional accuracy fell within the range of 0.57–0.87. Depending on the complexity of the markets and availability of data, the directional accuracy of financial time series typically ranges between 0.55 and 0.80 (Bustos and Pomares-Quimbaya,, 2020).
@@
When a movement prediction is better than a random guess, it also indicates that the input data contain useful information that has been leveraged effectively to predict future movements. Borrowing the "less is more" principle from cognitive science, where reducing information complexity can lead to an improved cognitive process (Gigerenzer and Brighton,, 2009), we can subsequently use the result of the movement prediction to generate the point forecast, using only the binary labels as an additional predictor along with the original naïve forecast, as shown in Figure 1. The binary labels can then be regarded as "verified predictors" for the target variable, and no more noise is introduced to affect the final point forecast. Unfortunately, the use of movement prediction results only to improve point forecasts remains unexplored in the forecasting community. Most studies either stopped at the movement prediction stage or used a direct regression approach to conduct point forecasting. Although it is often argued that, in many cases, compared with point forecasting, movement prediction is more important for decision-making (Sezer et al.,, 2020), the final decision generally still depends on the estimated value of the target variable. For example, in financial markets, the decision could be to trade at a specific price given a predicted future movement. Therefore, the decision maker actually conducted the classification-to-regression conversion, albeit implicitly. However, how and why movement prediction results can be systematically exploited to improve point forecasting in random walk-like series is largely unknown.
@@
Under these conditions, this study aimed to explore the potential of using movement prediction to enhance point forecasting in random walk-like time series. In particular, the target variable is assumed to follow a symmetric random walk process, which is the basic type of random walk characterized by the absence of drift and equal probabilities of upward and downward movements (Ibe,, 2014). By reformulating the symmetric random walk process to incorporate the future movement of the target variable, the point forecast can naturally be expressed as the original naïve forecast adjusted by movement prediction, which is referred to as the movement-prediction-adjusted naïve forecast. The weight applied to the movement term is determined by two parameters: one reflecting the directional accuracy and the other representing the mean absolute increment. The setting of these two parameters introduces an inherent trade-off and uncertainty into the prediction performance. Specifically, larger parameter values suggest potentially meaningful improvements over the original naïve forecast, whereas smaller parameter values indicate a greater likelihood that the adjusted naïve forecast will outperform the original naïve forecast. To balance this trade-off, the parameters can be empirically determined by sliding windows on the in-sample data.
@@
Since, in practice, movement prediction can exhibit a wide range of directional accuracies, the adjusted naïve forecast should be evaluated across different accuracy levels to examine the impact of directional accuracy on the performance of the proposed method. Therefore, we use synthetic time series to simulate multiple forecasting scenarios, assuming that for each directional accuracy level, diverse movement predictions are provided. For all the experiments, the evaluation metrics include the RMSE, mean absolute error (MAE), mean absolute percentage error (MAPE), and symmetric mean absolute percentage error (sMAPE), and the original naïve forecast serves as the sole baseline. In this experimental setting, it can be determined whether the improvement over the baseline by the proposed method is statistically significant.
This study makes three significant contributions to the literature. First, this study provides new insights into the predictability of random walk-like series, addressing a longstanding challenge in the field. When meaningful movement prediction was provided, the point values of the random walk-like series became partially predictable. Second, by illuminating the relationship between the actual and estimated parameter values, this study underscores the source of uncertainty in the effectiveness of the proposed method, a critical consideration often underemphasized by purely data-driven approaches. Third, the proposed method bridges movement prediction and point forecasting, which are two distinct tasks in time series forecasting. Although the use of binary classification to guide regression is well established, relying solely on the movement prediction of the target variable for its own point forecasting offers a fresh perspective.
@@
The remainder of this paper is organized as follows. Section 2 introduces the details of the movement-prediction-adjusted naïve forecast, explains how to derive it, and explains why it can outperform the original naïve forecast. Section 3 presents a simulation study that assesses the performance of movement-prediction-adjusted naïve forecasts using synthetic datasets in a controlled experimental setting. Section 4 discusses the key findings, implications, and limitations of the study. Finally, Section 5 concludes the paper and outlines future research directions.
2 Movement-Prediction-Adjusted Naïve Forecast
@@
This section presents the details of the movement-prediction-adjusted naïve forecast for time series that exhibit symmetric random walk characteristics. We first derive the general expression of the adjusted naïve forecast based on the definition of symmetric random walks, then explain how to optimize the parameters attached to the movement term, and finally, how to adjust the parameters to improve the forecasting effectiveness and detail the procedure for generating forecasts on the out-of-sample set.
2.1 General Expression
@@
where ${y}_{t}$ is the point value at time step $t$ and ${y}_{t-1}$ is the point value at time step $t-1$. For random walks exhibiting homoscedasticity, the variance $\sigma_{t}^{2}$ remains constant over time, whereas for those exhibiting heteroscedasticity, the variance $\sigma_{t}^{2}$ varies over time. The symmetric nature of the random increment ${{\epsilon}_{t}}$ with a mean of 0 ensures that there is no inherent bias in the movement direction of the target variable and that upward and downward movements are equally likely (Ibe,, 2014). Furthermore, the random increment $\epsilon_{t}$ can be expressed as a product of two components: a nonnegative random increase $|\epsilon_{t}|$ representing the magnitude of the increment and a sign component ${D}_{t}$, which takes a value of $+1$ for upward movement and $-1$ for downward movement, indicating the direction of the increment between two adjacent points. This allows the process to be reformulated as follows:
@@
The nonnegative random increment $|\epsilon_{t}|$ can be further expressed as a product of two terms:
@@
where $\bar{\epsilon}$ represents the mean absolute increment calculated from the in-sample data, which can be interpreted as a base magnitude of the increments. The term $\theta_{t}$ is introduced as a nonnegative parameter that controls the magnitude of the increment at time step $t$. Eq. (3) implies that the point forecast of the target variable can be generated if ${{D}_{t}}$ can be predicted and both $\theta_{t}$ and $\bar{\epsilon}$ can be estimated. Consider a simple scenario in which the magnitude of the increment between two adjacent points is fixed for future time steps; then, $\theta_{t}$ can be set to a constant value. Consequently, the point forecast for future time step $t$, denoted as $\hat{y}_{t}$, can be expressed as follows:
@@
where ${{\hat{D}}_{t}}$ is the movement prediction, $\theta$ is a fixed scalar, and $\theta\in\mathbb{Q}$. Therefore, the point forecast generated via Eq. (4) can be regarded as an adjustment of the original naïve forecast through a weighted movement prediction term; hence, it is named the movement-prediction-adjusted naïve forecast.
2.2 Movement Term Weight
According to Eq. (4), the weight of the movement term in the adjusted naïve forecast is the product of two parameters, $\theta$ and $\bar{\epsilon}$, both of which can be estimated from the in-sample data. Since $\bar{\epsilon}$ is directly computed as the mean absolute increment of the in-sample series, only $\theta$ requires optimization. This is typically achieved by minimizing the mean squared error (MSE) of the adjusted naïve forecast for the in-sample set. Because the MSE of the original naïve forecast is a fixed value on the in-sample set, minimizing the MSE of the adjusted naïve forecast is mathematically equivalent to maximizing the difference between the MSEs of the original and adjusted naïve forecasts:

### v8

@@ Abstract
This study introduces a movement-prediction-adjusted naïve forecast for time series exhibiting symmetric random walk characteristics, which is applicable after accurate movement predictions are available. Specifically, the original naïve forecast is adjusted by a weighted movement prediction term, where the weights are determined via two parameters derived from the in-sample data: one based on directional accuracy of the movement prediction and the other on the mean absolute increment of the target series. Simulation experiments were conducted across four types of synthetic symmetric random walk series, each with different variance structures. For each time series, diverse movement predictions with predefined directional accuracies were randomly generated, and the resulting forecasts were evaluated via the RMSE, MAE, MAPE, and sMAPE metrics. The results demonstrated a clear monotonic improvement in the forecast performance as the directional accuracy increased. Notably, the adjusted naïve forecast achieved statistically significant improvements even at relatively low directional accuracy levels slightly above 0.50. These findings imply that the movement-prediction-adjusted naïve forecast can serve as an effective second-stage method for forecasting symmetric random walk time series when consistent and accurate movement predictions are provided.
@@ 1 Introduction
Time series forecasting plays a crucial role in many decision-making processes. The widespread demand for accurate point forecasts across various domains has attracted considerable attention from researchers and practitioners. However, not all time series are equally predictable. Time series exhibiting random walk characteristics pose significant forecasting challenges (Hamilton,, 2020). In particular, financial time series, known for these characteristics and their practical importance in financial markets, have been extensively studied as test cases for forecasting theories and methods. One prominent theory is the efficient market hypothesis (EMH), which suggests that asset prices incorporate all available information, rendering historical prices unpredictable (Fama,, 1970). For a random walk-like series, the naïve forecast, which assumes that the future value at the next time step is equal to the current value, is often regarded as the optimal point forecast.
@@
In support of the EMH, Moosa, (2013) reported that as the volatility of financial time series increases, the root mean square error (RMSE) of the forecasting models increases more quickly than that of the naïve forecast. Moreover, Moosa and Burns, (2014) argued that claims of outperforming the naïve forecast in terms of the RMSE are often misleading, frequently involving introduced dynamics without rigorous statistical validation. Despite numerous attempts to exploit weak structural signals in financial time series via various predictors, existing forecasting methods—including statistical, machine learning, and deep learning approaches—have typically failed to outperform naïve forecasts (Kilian and Taylor,, 2003; Moosa and Burns,, 2016; Thakkar and Chaudhari,, 2021; Petropoulos et al.,, 2022; Ellwanger and Snudden,, 2023; Hewamalage et al.,, 2023; Zeng et al.,, 2023; Beck et al.,, 2025).
@@
From the forecaster’s perspective, the unpredictability of a variable could be attributed to incomplete knowledge of the underlying processes controlling the data, ignorance of the various contributing factors, and imperfections in analytical methods (Laplace,, 2012). In other words, randomness in data is not inherent but rather apparent (Crutchfield and Feldman,, 2003). Although useful information may be contained in the input space, it can be overwhelmed by irrelevant data, making it difficult for current regression methods to exploit such information (John et al.,, 1994; Guyon and Elisseeff,, 2003; Verleysen and François,, 2005). Moreover, identifying the predictors on which the target variable truly depends is a complex task (Granger,, 1969; Stock and Watson,, 2002). However, these theories and empirical evidence reveal only the practical difficulty of finding exploitable patterns; they do not eliminate the possibility of discovering any predictability in the data.
@@
Given the difficulty of generating precise numerical predictions from noisy data in random walk-like time series, researchers have increasingly shifted their focus from point forecasting to movement prediction, which involves predicting whether the value will increase or decrease at the next time step. Consequently, the regression task was reduced to a binary classification task. For a pure random walk, the probability that future movement will increase or decrease is 0.5, and the mutual information between the predictors and the target is zero (Kraskov et al.,, 2004). For a random walk-like time series using abundant predictors, the mutual information can be greater than zero, indicating that there exists some dependency between the predictors and the target (Cover and Thomas,, 2012). Practically, this means that the predictors reduce the uncertainty of the binary outcome. Although this does not guarantee high predictive accuracy, it suggests that, theoretically, one can build a model that performs better than random guessing (Ircio et al.,, 2020). In contrast, for point forecasting in a random walk-like series, the continuous nature of the target means that even if a small gain in mutual information exists, it only marginally reduces the overall uncertainty and does not sufficiently narrow down the precise numerical result (Beraha et al.,, 2019).
@@
Indeed, movement prediction is often considered easier than point forecasting in financial time series (Taylor,, 2008; Moosa,, 2013). Studies in the field of behavioral finance suggest that investor sentiment and psychology can lead to certain patterns (De Bondt and Thaler,, 1985; Lee et al.,, 1991; Baker and Wurgler,, 2007; Barberis et al.,, 1998), which could lead to mutual information between the sentiment predictors and the binary target being greater than zero and making future movements predictable. Therefore, sentiment indicators are often used to improve the directional accuracy in these movement prediction tasks. For example, Weng et al., (2017) incorporated Google news counts and Wikipedia page views alongside historical data to predict the next day’s movement of AAPL stock via support vector machines (SVMs), achieving up to 0.80 directional accuracy. Ma et al., (2023) proposed a multisource aggregated classification (MAC) method for the prediction of stock price movement, incorporating the numerical features and market-driven news sentiments of target stocks, as well as the news sentiments of their related stocks, and the directional accuracy can reach as high as 0.70. Wang et al., (2025) used multimodal data that included news data to train a hybrid model consisting of a long short memory network (LSTM) and transformer for the prediction of stock movement, and the directional accuracy fell within the range of 0.57–0.87. Depending on the complexity of the markets and availability of data, the directional accuracy of financial time series typically ranges between 0.55 and 0.80 (Bustos and Pomares-Quimbaya,, 2020).
@@
When a movement prediction is better than a random guess, it also indicates that the input data contain useful information that has been leveraged effectively to predict future movements. Inspired by the ’less is more’ principle, which suggests that reducing informational complexity can improve decision-making (Gigerenzer and Brighton,, 2009), movement prediction can further serve as a simplified predictor in addition to the naïve forecast to generate the final point forecast, as shown in Figure 1. This indirect approach requires only the naïve forecast and the movement prediction as predictors, eliminating the need for additional noisy predictors. The naïve forecast, which represents the current value, is most closely related to the future value, whereas the movement prediction contributes new directional information not contained within the naïve forecast. Unfortunately, the use of movement prediction alone to improve naïve forecasts remains unexplored in the forecasting community. Previous studies either perform only movement prediction or apply regression directly to noisy data. Although it is often argued that, in many cases, compared with point forecasting, movement prediction is more important for decision-making (Sezer et al.,, 2020), the final decision generally still depends on the estimated value of the target variable. For example, in financial markets, the decision could be to trade at a specific price given a predicted future movement. Thus, even when practitioners rely primarily on predicted movements, they implicitly perform a classification-to-regression conversion by translating the predicted directions into quantitative trading decisions. However, there remains a notable lack of systematic approaches for integrating movement prediction outcomes with naïve forecasts in the context of forecasting random walk-like time series.
@@
Under these conditions, this study aimed to explore the potential of using movement prediction to enhance point forecasting in random walk-like time series. To this end, we propose a second-stage method that effectively integrates movement prediction with the naïve forecast after accurate and consistent movement predictions have been obtained. The target variable is assumed to follow a symmetric random walk process, which is the basic type of random walk characterized by the absence of drift and equal probabilities of upward and downward movements (Ibe,, 2014). By reformulating the symmetric random walk process to incorporate the future movement of the target variable, the point forecast can naturally be expressed as the original naïve forecast adjusted by the movement prediction term. Hence, it is named the movement-prediction-adjusted naïve forecast. The weight applied to the movement term is determined by two parameters derived from the in-sample data: one based on the directional accuracy of the movement prediction and the other on the mean absolute increment of the target series. The setting of these two parameters involves an inherent trade-off on the prediction performance. While larger parameter values indicate a lower likelihood of achieving improvements through adjustment, these improvements can be substantial; smaller values increase the likelihood of improvement, although such improvements may be marginal or negligible compared with the naïve forecast. To balance this trade-off, the parameters can be empirically estimated by sliding windows on the in-sample data.
@@
Since, in practice, movement prediction can exhibit a wide range of directional accuracies, the adjusted naïve forecast should be evaluated across different accuracy levels to examine the impact of directional accuracy on the performance of the proposed method. Therefore, we use synthetic time series to simulate multiple forecasting scenarios, assuming that for each directional accuracy level, diverse movement predictions are provided. For all the experiments, the evaluation metrics include the RMSE, mean absolute error (MAE), mean absolute percentage error (MAPE), and symmetric mean absolute percentage error (sMAPE), and the original naïve forecast serves as the sole baseline. In this experimental setting, we can evaluate whether the improvement over the baseline by the proposed method is statistically significant.
This study makes three significant contributions to the literature:
  • First, this study offers new insights into the predictability of random walk-like series, presenting a feasible approach to address this longstanding challenge in the field. When meaningful movement prediction is available, the point values of these series, which are traditionally viewed as unpredictable, exhibit conditional predictability.
  • Second, the proposed method bridges movement prediction and point forecasting, which are two distinct tasks in time series forecasting. Although the use of binary classification to guide regression is well established, relying solely on the movement prediction of the target variable for its own point forecasting offers a fresh perspective.
  • Third, this study introduces a clear boundary criterion, which explicitly states that the adjusted naïve forecast outperforms the original naïve forecast when the product of the actual optimal parameter values exceeds half the product of the estimated optimal parameter values. This contribution advances forecasting methodologies by clarifying the conditions required for improved performance.
@@
The remainder of this paper is organized as follows. Section 2 introduces the details of the movement-prediction-adjusted naïve forecast, explains how to derive it and under what conditions it can outperform the original naïve forecast. Section 3 presents a simulation study that assesses the performance of movement-prediction-adjusted naïve forecasts via synthetic datasets in a controlled experimental setting. Section 4 discusses the key findings, implications, and limitations of the study. Finally, Section 5 concludes the paper and outlines future research directions.
2 Method
@@
This section presents the details of the movement-prediction-adjusted naïve forecast. We first derive the general expression of the adjusted naïve forecast based on the definition of symmetric random walks and then explain how to optimize the parameters to improve the forecasting effectiveness and detail the procedure for generating forecasts on the out-of-sample set.
2.1 General Formulation of the Movement-Prediction-Adjusted Naïve Forecast
@@
where ${y}_{t}$ is the point value at time step $t$ and ${y}_{t-1}$ is the point value at time step $t-1$. For random walks exhibiting homoscedasticity, the variance $\sigma_{t}^{2}$ remains constant over time, whereas for those exhibiting heteroscedasticity, the variance $\sigma_{t}^{2}$ varies over time. The symmetric nature of the random increment ${{\epsilon}_{t}}$ ensures that there is no inherent bias in the movement direction of the target variable and that upward and downward movements are equally likely (Ibe,, 2014). Furthermore, the random increment $\epsilon_{t}$ can be expressed as a product of two components: a nonnegative term $|\epsilon_{t}|$ representing the magnitude of the increment and a sign component ${D}_{t}$, which takes a value of $+1$ for upward movement and $-1$ for downward movement, indicating the direction of the increment between two adjacent points. This allows the process to be reformulated as follows:
@@
The nonnegative term $|\epsilon_{t}|$ can be further expressed as a product of two terms:
@@
where $\bar{\epsilon}$ represents the mean absolute increment calculated based on the in-sample data, which can be interpreted as a base magnitude of the increments. The term $\theta_{t}$ is introduced as a nonnegative parameter that controls the magnitude of the increment at time step $t$. Eq. (3) implies that the point forecast of the target variable can be generated if ${{D}_{t}}$ can be predicted and both $\theta_{t}$ and $\bar{\epsilon}$ can be estimated. Consider a simplified scenario in which the magnitude of the increment between two adjacent points is fixed for future time steps; in this case, $\theta_{t}$ can be set to a constant value. Consequently, the point forecast for future time step $t$, denoted as $\hat{y}_{t}$, can be expressed as follows:
@@
where ${{\hat{D}}_{t}}$ is the movement prediction, $\theta$ is a fixed scalar, and $\theta\in\mathbb{Q}$. Suppose that the movement prediction has been provided, the point forecast generated via Eq. (4) can be regarded as an adjustment of the original naïve forecast through a weighted movement prediction term; hence, it is named the movement-prediction-adjusted naïve forecast.
2.2 Parameter Optimization Using In-Sample Data
According to Eq. (4), the weight of the movement term in the adjusted naïve forecast is the product of two parameters, $\theta$ and $\bar{\epsilon}$, both of which can be optimized based on the in-sample data. Because $\bar{\epsilon}$ is computed directly as the mean absolute increment of the in-sample series, only $\theta$ requires optimization. This optimization is typically achieved by minimizing the mean squared error (MSE) of the adjusted naïve forecast within the in-sample set. Given that the original naïve forecast’s MSE is constant for the in-sample set, minimizing the adjusted naïve forecast’s MSE is equivalent to maximizing their difference: